6. A 2.00-kg rock is released from rest at a heigh
Answer #1

a) KE at h = 30 m

KE = 0

b) PE at H = 30 m

PE = mgh = 2*9.8*30 = 588 J

c) TE = PE + KE = 0 + 588 = 588 J

d) h= 0

KE = 0.5*mv^2

v^2 = 2gh

KE = 0.5*2*2gh = 588 J

e) PE at h=0

PE = mgh = 0

f) TE +PE + KE = 588 J

7) s = ut + 0.5*at^2

0.62 = 0 + 0.5*a*1^2

a = 1.24 m/s^2

F = ma = 85*1.24 = 105.4 N
Answer #2

a= 0.62-0/1 = 0.62

f= ma= 85*0.62=52.7N